# Steady-states and response functions of the periodically driven O($N$) scalar field theory

### v1

@@ I Introduction
The past decade has seen remarkable progress in realizing new light-induced phases of matter. Examples include superconductivity in K3C60 [Mitrano16; Dodge23; Comment23; Dodge25], the cuprates [Cremin19; Shimano23; Shimano24; Cavalleri24; Comin25; Michael2024; Michael2025], and the nickelates [Xu25], charge density waves [Gedik20; Gedik21; Carpene24; Gedik25a; Gedik25b; Padma2025], antiferromagnetism [Gedik24] and ferroelectricity [Nova2019; Fechner2024]. Motivated by these results, we examine a general theory of order parameter dynamics in the presence of a time-periodic modulation of the coupling constants of the effective free energy of the order parameter. We assume that the periodic driving does not break any symmetry associated with the order parameter, and study the phase diagram in which the symmetry may be spontaneously broken in some phases. Our focus is on the regime where the natural frequency of order parameter dynamics is comparable to the driving frequency $\Omega$, and the damping from the environment is moderately weak. In such a regime, a Floquet point-of-view is appropriate at linear order, where we consider frequencies shifted by integer multiples of $\Omega$. However, we will also study the full non-linear case, where all frequencies couple.
Figure 1: Panel (a): Phase diagram obtained by solving Eqs. (4) with parameters $u=\Omega^{2}$, and $\gamma=0.03\,\Omega$. Six phases appear: normal (gray), superconducting (SC,blue), period-doubling SC (red), pair density wave (PDW, blue hatches), PDW with period doubling (red hatches) and a quasi-periodic PDW (yellow). The symbols mark the points whose order-parameter time traces over the last ten drive periods are shown in panel (b). As detailed in App. D, we also performed a dynamical stability analysis. The hatched regions indicate an instability of the normal state towards a PDW, oscillating at either half the drive frequency (red hatched regions), labeled as PDW with period doubling (PDW w PD), or a at the drive frequency, labeled as PDW withoud period doubling (PDW w/o PD). The yellow region indicates a portion of the phase diagram where the steady state is non-periodic but the correlation function $C_{\mathbf{q}}(t)$ diverges at finite $\mathbf{q}$. We labeled this region as quasi-periodic PDW (qp-PDW). Panel (b): Representative dynamics $\bar{b}(t)$ over the last ten periods, plotted versus $t/T_{0}$, for the parameter points indicated by the symbols in panel (a). Top: metal ($\bar{b}=0$), middle: SC (small oscillations with the drive period about a nonzero mean), bottom: period-doubling SC (oscillations with period $2T_{0}$ around zero).
@@
so that $r_{0}$ is the tuning parameter across the equilibrium transition, and $r_{1}$ is the amplitude of the periodic driving. A number of earlier studies have obtained related results on similar models [Chandran16; Diehl16; Millis20; DaiLee21; Mitra21a; Mitra21b; Dai22; Sols24; Chandra25a; Moessner25; Zelle25; ZelleMillis25; Chandra25b]. We go beyond these works and obtain the phase diagram in the presence of both self-interaction and dissipation. The phase diagram is shown in Fig. 1a as a function $r_{0}/\Omega^{2}$ and $r_{1}/\Omega^{2}$, and is obtained by various analytic arguments, and a full numerical solution of the large $N$ equations. We use the terminology of the $N=2$ case which applies to the superconductor-normal transition. The long-time steady-state time-dependence of the O($N$) order parameter (SC) $b$ is shown in Fig. 1b. In the ‘SC’ phase, the order parameter oscillates at a frequency $\Omega$ and non-zero mean, while in the ‘SC with PD’ phase (sometimes called a ‘time crystal’) there is a period-doubling and the order parameter oscillates at a frequency $\Omega/2$ with zero mean. Our dynamical stability analysis also unveils phases were the order parameter condenses at finite momentum, similar to previous discussions in the context of parametrically driven bosonic system [Dai2021; Chandra25a; Chandra25b; Kaplan2025b].
Section V obtains results on response functions of the various phases of Fig. 1. For the $N=2$ case, the response function is the optical conductivity, associated with coupling of an external electromagnetic field to the conserved O($N$) current. We show that whenever the order parameter develops an instantaneous expectation value, a mass term for the photon appears. When the order parameter oscillates around a large offset, this leads to a conventional Meissner effect, with the magnetic field being fully expelled from the material. Conversely, when such offset is small or vanishing, a fraction of the external magnetic field can penetrate the sample in the form of a standing wave, a Meissner polariton, that is, a collective mode emerging from coupling the order parameter oscillations to light.
@@ II Model and large-$N$ limit
$\mathcal{L}[b]=\frac{1}{2}(\partial_{\mu}b)^{2}-\frac{r}{2}b^{2}-\frac{u}{4N}b^{4}\,$, (2)
where $b^{2}\equiv\sum_{\alpha}b_{\alpha}^{2}$, and $u>0$. We assume then that the parameter $r$, which in equilibrium describes the distance from the critical point, is periodically driven as in Eq. (1). Model (2) can describe the dynamics of several order parameters that are not associated with a conserved quantity, such as an antiferromagnetic order parameter for $N=3$. Even though a superconducting order parameter is usually the conjugate variable of a conserved quantity, that is the particle number, model (2) can also be employed for $N=2$ to describe the dynamics of the superconducting order parameter in the Bose-Hubbard model [Dai22] or in the attractive Fermi-Hubbard model near integer filling, where a particle-hole symmetry emerges. Furthermore, model (2) can be thought of as an effective theory for the Goldstone modes in a superfluid [Bogoliubov1947; Anderson1958] or in a superconductor, provided that the electromagnetic field is regarded as non-dynamical [AmuNote].
@@
In absence of a drive and for the $N=2$, model (3) describes the dynamics of the metal-to-superconducting transition [Tinkham2004]. In this case, for $r\geq r_{c}$ the order parameter expectation value is zero, corresponding to a symmetric, normal phase, while for $r<r_{c}$, the order parameter acquires a finite expectation value corresponding to the symmetry-broken, superconducting phase.
Eq. (3) can be solved analytically only in the large-$N$ limit. To this end, one performs a Hubbard-Stratonovich transformation to decouple the quartic interaction term and then derive the saddle-point equations in the large-$N$ limit. The derivation is shown in App. A, and we report here only the final result. We assume a spatially homogeneous condensate and we posit $\langle b\rangle=(\bar{b},0,\dots,0)$. The equation of motion for the order parameter and the correlation functions $C_{\mathbf{q}}(t)=\langle b(\mathbf{q},t)^{2}\rangle-\bar{b}^{2}$ and $D_{\mathbf{q}}(t)=\langle[\partial_{t}b(\mathbf{q},t)]^{2}\rangle-(\partial_{t}\bar{b})^{2}$ are given by:
@@
$\lambda(t)=r(t)+u\left[\bar{b}^{2}+\int_{\mathbf{q}}C_{\mathbf{q}}(t)\right],$ (5)
@@
We notice that similar equations were derived for an isolated quantum system ($\gamma=0$) in Refs [Chandran2016; Natseh2021_2]. In this respect, here we extend their analysis to an open system and show that multiple steady states can be stabilized, depending on the strength of $r_{1}$.
@@ III Floquet Analysis
Given the presence of a periodic drive and dissipation, we expect the system to reach a periodic steady state, i.e., a state where all observables are periodic, although not necessarily with the same period as the drive, as discussed further below. By assuming that, in the periodic steady state, $\lambda(t)$ oscillates with the same period as the drive, we observe that Eq. (4a) takes the form of a Mathieu-Hill equation [Mathieu; Hill]. In this case, nontrivial solutions can be found using Floquet’s theorem [Floquet], i.e., by working in the Fourier basis spanned by integer multiples of the drive frequency. The order parameter can then be written as
@@
where $\sum_{n}\equiv\sum_{n=-\infty}^{+\infty}$, and $\nu\in(-\Omega/2,\Omega/2]$ is a parameter allowing for a different periodicity than the drive itself. Inserting (6) into (4a) and expanding $\lambda(t)=\sum_{n}\lambda_{n}e^{-in\Omega t}$, we get
@@
Figure 2: Lines in $\lambda_{0}$-$\lambda_{1}$ plane where the damped Mathieu equation supports oscillating steady states at $\omega=0$ (blue lines) and at $\omega=\Omega/2$ (red lines) for $\gamma=0.3\Omega$. In the gray shaded regions, the solutions are decaying to zero, whereas in the unshaded regions solutions grow indefinitely with time.
In Fig. 2 we plot the lines in the $\lambda_{0}$-$|\lambda_{1}|$ plane where $\mathrm{det}\,\mathcal{K}(\nu)=0$ for $\gamma/\Omega=0.3$ (solid blue and red lines). For any finite $\gamma$, $\mathcal{K}(\nu)$ has a vanishing eigenvalue only for $\nu=0,\Omega/2$.
The blue (red) lines in Fig. 2 indicate the lines in parameter space where the Mathieu equation supports oscillating steady states with $\nu=0$ ($\nu=\Omega/2$). The gray area indicates the parameter space where only solutions that decay to zero are supported. Conversely, in the white regions a parametric resonance occurs and $b(t)$ grows indefinitely with time. These regions are known as Arnold tongues in the mathematical literature [Arnold1965; Nayfeh1979; Kovacic2018; Strogatz2018].
@@ III.1.3 Discussion of phase diagram and excitations
1st order PT and exceptional points — We then notice that the order parameter grows continuously from zero to a finite value across the solid black line, and discontinuously across the dashed black: This suggest that the phase transition is of the second order across the solid line, and of the first order across the dashed line. This peculiar behavior of the order parameter is typical of driven-dissipative systems which feature an exceptional point (EP) in the non-Hermitian matrix determining the order parameter [hanai2019non; hanai2020critical]. An EP is a point where two (or more) eigenstates that characterize the dynamics coalesce owing to the non-Hermitian nature of the system, such that they lose their completeness, leading to a spectral singularity. In the present case, the matrix $\mathcal{K}(\Omega/2)$ shows an exceptional point at $r_{0}=\Omega^{2}/4$, $r_{1}=\gamma\Omega/2$, i.e., where the black and the dashed lines join.
@@
Excitation in the normal phase— We start from the disordered normal phase. To gain insight on the structure of the excitations, we first consider the case with $\lambda_{1}=0$. In this case the retarded Green’s function describes a system of two decoupled modes (infinite decoupled modes in the case without truncation). Each mode has a complex energy with two branches (as for a damped harmonic oscillator). The real part of the energy is centered around either 0 or $\Omega$ (in general around $n\Omega$ for the $n$-th mode, for the non truncated case), and they cross at $q^{*}=\sqrt{-r_{0}+(\Omega^{2}+\gamma^{2})/4}$, provided that $r_{0}<(\Omega^{2}+\gamma^{2})/4$. The imaginary part of the energy (i.e., the inverse lifetime), instead, is the same for all modes. Both real and imaginary parts show a bifurcation, which is typical of driven-dissipative non-Hermitian systems [heiss2012physics]. A finite value of the intermode coupling $\lambda_{1}$ creates a non-Hermitian avoided crossing of the frequencies (i.e., bifurcations) close to the momenta $q^{*}$ This emerges as a softening in the imaginary part of the energy around $q^{*}$ (see panel (b) bottom in Fig. 4)
@@
By approaching the critical curves, the imaginary parts of the dispersions further softens until they become zero for some momenta, signaling the emergence of an ordered phase. The softening appears different depending on which side is approached: Upon approaching the second-order phase transition line (panel c bottom in Fig. 4), the softening happens at $q=0$, while approaching the first-order phase transition line (panel b bottom in Fig. 4 ), the finite momentum $q^{*}$ softens. This is remarkable, as this would signal an instability at finite momenta, while we assumed an ordered translation-invariant phase. This suggests that our homogeneous Ansatz for the ordered phase is incomplete, and the inclusion of a spatially-modulated order parameter Ansatz would in fact show the emergence of a phase with broken translational symmetry near the first order transition line. The emergence of spatially patterned phases in driven systems is known and correspond to Faraday waves, which are observed, e.g., in periodically driven condensates, see, e.g., [Faraday1; Faraday2; Faraday3]. However, this goes beyond the scope of our work and we will leave it for further investigations.
@@
Excitations in the superconducting phase— We consider now the fluctuations in the ordered superconducting phase. In this case there is always a mode with vanishing imaginary and real part at $q=0$, which corresponds to the Goldstone mode of the ordered phase (panel a, right figure, in Fig. 4). The mode is purely diffusive as $\text{Im}(E(q))\propto-q^{2}$ for small $q$. This is well known for driven-dissipative systems and has been predicted and experimentally verified, e.g., for exciton-polariton systems [Wouters2007; claude2025observation]. Note that such mode is a manifestation of Floquet-Nambu-Goldstone modes discussed in Ref. [Sols24].

### v2

@@ I Introduction
The past decade has seen remarkable progress in realizing possible new light-induced phases of matter. Examples include superconductivity in K3C60 [44], the cuprates [9; 52; 53; 19; 30; 43; 42], and the nickelates [65], charge density waves [36; 68; 56; 60; 51; 55], antiferromagnetism [28] and ferroelectricity [54; 20] (we note that the experimental observation of light-induced superconducting phases has been recently challenged [16; 5; 17]). Motivated by these results, we examine a general theory of order parameter dynamics in the presence of a time-periodic modulation of the coupling constants of the effective free energy of the order parameter. We assume that the periodic driving does not break any symmetry associated with the order parameter, and study the phase diagram in which the symmetry may be spontaneously broken in some phases. Our focus is on the regime where the natural frequency of order parameter dynamics is comparable to the driving frequency $\Omega$, and the damping from the environment is moderately weak. In such a regime, a Floquet point-of-view is appropriate at linear order, where we consider frequencies shifted by integer multiples of $\Omega$. However, we will also study the full non-linear case, where all frequencies couple.
Figure 1: Panel (a): Phase diagram obtained by solving Eqs. (4) with parameters $u=\Omega^{2}$, and $\gamma=0.03\,\Omega$. Six phases appear: normal (gray), superconducting (SC,blue), period-doubling SC (red), pair density wave (PDW, blue hatches), PDW with period doubling (red hatches) and a quasi-periodic PDW (yellow). The symbols mark the points whose order-parameter time traces over the last ten drive periods are shown in panel (b): the circle at $(r_{0}-r_{0,c})/\Omega^{2}=-0.4$, $r_{1}/\Omega^{2}=0.1$, the square at $(0.6,0.34)$, and the diamond at $(1.25,0.1)$. As detailed in App. D, we also performed a dynamical stability analysis. The hatched regions indicate an instability of the normal state towards a PDW, oscillating at either half the drive frequency (red hatched regions), labeled as PDW with period doubling (PDW w PD), or a at the drive frequency, labeled as PDW without period doubling (PDW w/o PD). The yellow region indicates a portion of the phase diagram where the steady state is non-periodic but the correlation function $C_{\mathbf{q}}(t)$ diverges at finite $\mathbf{q}$. We labeled this region as quasi-periodic PDW (qp-PDW). Panel (b): Representative dynamics $\bar{b}(t)$ over the last ten periods, plotted versus $t/T_{0}$, for the parameter points indicated by the symbols in panel (a). Top: metal ($\bar{b}=0$), middle: SC (small oscillations with the drive period about a nonzero mean), bottom: period-doubling SC (oscillations with period $2T_{0}$ around zero).
@@
so that $r_{0}$ is the tuning parameter across the equilibrium transition, and $r_{1}$ is the amplitude of the periodic driving. A number of earlier studies have obtained related results on similar models [7; 61; 10; 46; 47; 12; 45; 32; 27; 33]. In particular, out-of-equilibrium, self-interacting O($N$) models have been considered in Refs. [63; 67; 14; 13; 66]. It was shown that they exhibit a rich phenomenology, displaying, among others, prethermal and limit cycle phases. In Ref. [26], resonant optical driving of a high-$T_{c}$ superconductor at the sum of its Higgs and Josephson plasma frequencies produces persistent subharmonic oscillations of the order parameter, identified as a Higgs time-crystalline phase.
In this work, we calculate the phase diagram in the presence of both self-interaction and dissipation. The phase diagram is shown in Fig. 1a as a function $r_{0}/\Omega^{2}$ and $r_{1}/\Omega^{2}$, and is obtained by various analytic arguments, and a full numerical solution of the large $N$ equations. We use the terminology of the $N=2$ case which applies to the superconductor-normal transition. The long-time steady-state time-dependence of the O($N$) order parameter (SC) $b$ is shown in Fig. 1b. In the ‘SC’ phase, the order parameter oscillates at a frequency $\Omega$ and non-zero mean, while in the ‘SC with PD’ phase (sometimes called a ‘time crystal’) there is a period-doubling and the order parameter oscillates at a frequency $\Omega/2$ with zero mean. Our dynamical stability analysis also unveils phases where the order parameter condenses at finite momentum, similar to previous discussions in the context of parametrically driven bosonic system [11; 32; 33; 31].
Section V obtains results on response functions of the various phases of Fig. 1. For the $N=2$ case, the response function is the optical conductivity, associated with coupling of an external electromagnetic field to the conserved O($N$) current. We show that whenever the order parameter develops an instantaneous expectation value, a mass term for the photon appears. When the order parameter oscillates around a large offset, this leads to a conventional Meissner effect, with the magnetic field being fully expelled from the material. Conversely, when such offset is small or vanishing, a fraction of the external magnetic field can penetrate the sample in the form of a standing wave, a Meissner polariton, that is, a collective mode emerging from coupling the order parameter oscillations to light. Moreover, we show that near the onset of a phase with an oscillating order parameter, the imaginary part of the optical conductivity, $\sigma_{2}(\omega)$, can be substantially enhanced and display a $1/\omega$ behavior over a broad interval of frequencies. This observation offers a possible explanation of the photo-induced enhancement of $\sigma_{2}(\omega)$ that has been experimentally observed. Previously, microscopic theories of driven electron-phonon systems have been proposed as a possible explanation of light-induced superconductivity [37; 35; 3; 34; 18].
@@ II Model and large-$N$ limit
[…]
where $b^{2}\equiv\sum_{\alpha}b_{\alpha}^{2}$, and $u>0$. We assume then that the parameter $r$, which in equilibrium describes the distance from the critical point, is periodically driven as in Eq. (1). Model (2) can describe the dynamics of several order parameters that are not associated with a conserved quantity, such as an antiferromagnetic order parameter for $N=3$. Even though a superconducting order parameter is usually the conjugate variable of a conserved quantity, that is the particle number, model (2) can also be employed for $N=2$ to describe the dynamics of the superconducting order parameter in the Bose-Hubbard model [12] or in the attractive Fermi-Hubbard model near integer filling, where a particle-hole symmetry emerges.
@@
In absence of a drive and for the $N=2$, model (3) describes the dynamics of the metal-to-superconducting transition [62]. In this case, for $r\geq r_{c}$ the order parameter expectation value is zero, corresponding to a symmetric, normal phase, while for $r<r_{c}$, the order parameter acquires a finite expectation value corresponding to the symmetry-broken, superconducting phase.
Eq. (3) can be solved analytically only in the large-$N$ limit. To this end, one performs a Hubbard-Stratonovich transformation to decouple the quartic interaction term and then derive the saddle-point equations in the large-$N$ limit. It is known [4] that two-loop processes, which are excluded in this limit, play an important role in redistributing momentum and energy, which will eventually help the system thermalize. However, we do not aim to describe intrinsic many-body thermalization here, but rather the emergence of dynamical instabilities and non-equilibrium steady states in a driven open system. As it will become clear in what follows, a large-$N$ approximation is sufficient to obtain these phenomena. Moreover, we model damping and relaxation effects phenomenologically by introducing a Markovian noise.
The derivation of the large-$N$ equations is shown in App. A, and we report here only the final result. We assume a spatially homogeneous condensate and we posit $\langle b\rangle=(\bar{b},0,\dots,0)$. The equation of motion for the order parameter and the correlation functions $C_{\mathbf{q}}(t)=\langle b(\mathbf{q},t)^{2}\rangle-\bar{b}^{2}$ and $D_{\mathbf{q}}(t)=\langle[\partial_{t}b(\mathbf{q},t)]^{2}\rangle-(\partial_{t}\bar{b})^{2}$ are given by:
@@
$\lambda(t)=r(t)+u\left[\bar{b}^{2}(t)+\int_{\mathbf{q}}C_{\mathbf{q}}(t)\right],$ (5)
@@
We notice that similar equations were derived for an isolated quantum system ($\gamma=0$) in Refs [6; 48]. In this respect, here we extend their analysis to an open system and show that multiple steady states can be stabilized, depending on the strength of $r_{1}$.
@@ III Floquet Analysis
Given the presence of a periodic drive and dissipation, we expect the system to reach a periodic steady state, i.e., a state where all observables are periodic, although not necessarily with the same period as the drive, as discussed further below. By assuming that, in the periodic steady state, $\lambda(t)$ oscillates with the same period as the drive, we observe that Eq. (4a) takes the form of a Mathieu-Hill equation [41; 25]. In this case, nontrivial solutions can be found using Floquet’s theorem [21], i.e., by working in the Fourier basis spanned by integer multiples of the drive frequency. The order parameter can then be written as
@@
where $\sum_{n}\equiv\sum_{n=-\infty}^{+\infty}$, and $\nu\in(-\Omega/2,\Omega/2]$ is a parameter allowing for a different periodicity than the drive itself. As we shall see in Sec. IV, the assumption of $\lambda(t)$ being time periodic with period $2\pi/\Omega$ applies to the vast majority of the phase diagram in Fig. 1. Inserting (6) into (4a) and expanding $\lambda(t)=\sum_{n}\lambda_{n}e^{-in\Omega t}$, we get
@@
Figure 2: Lines in $r_{0}$-$|r_{1}|$ plane where the damped Mathieu equation supports oscillating steady states at $\omega=0$ (blue lines) and at $\omega=\Omega/2$ (red lines) for $\gamma=0.3\Omega$. In the gray shaded regions, the solutions are decaying to zero, whereas in the unshaded regions solutions grow exponentially with time.
In Fig. 2 we plot the lines in the $r_{0}$-$|r_{1}|$ plane where $\mathrm{det}\,\mathcal{K}(\nu)=0$ for $\gamma/\Omega=0.3$ (solid blue and red lines). For any finite $\gamma$, $\mathcal{K}(\nu)$ has a vanishing eigenvalue only for $\nu=0,\Omega/2$.
The blue (red) lines in Fig. 2 indicate the lines in parameter space where the Mathieu equation supports oscillating steady states with $\nu=0$ ($\nu=\Omega/2$). The gray area indicates the parameter space where only solutions that decay to zero are supported. Conversely, in the white regions a parametric resonance occurs and $b(t)$ grows indefinitely with time. These regions are known as Arnold tongues in the mathematical literature [2; 49; 38; 59].
@@ III.1.3 Discussion of phase diagram and excitations
1st order PT and exceptional points — We then notice that the order parameter grows continuously from zero to a finite value across the solid black line, and discontinuously across the dashed black: This suggest that the phase transition is of the second order across the solid line, and of the first order across the dashed line. This peculiar behavior of the order parameter is typical of driven-dissipative systems which feature an exceptional point (EP) in the non-Hermitian matrix determining the order parameter [22; 23]. An EP is a point where two (or more) eigenstates that characterize the dynamics coalesce owing to the non-Hermitian nature of the system, such that they lose their completeness, leading to a spectral singularity. In the present case, the matrix $\mathcal{K}(\Omega/2)$ shows an exceptional point at $r_{0}=\Omega^{2}/4$, $r_{1}=\gamma\Omega/2$, i.e., where the black and the dashed lines join.
@@
Excitation in the normal phase— We start from the disordered normal phase. To gain insight on the structure of the excitations, we first consider the case with $\lambda_{1}=0$. In this case the retarded Green’s function describes a system of two decoupled modes (infinite decoupled modes in the case without truncation). Each mode has a complex energy with two branches (as for a damped harmonic oscillator). The real part of the energy is centered around either 0 or $\Omega$ (in general around $n\Omega$ for the $n$-th mode, for the non truncated case), and they cross at $q^{*}=\sqrt{-r_{0}+(\Omega^{2}+\gamma^{2})/4}$, provided that $r_{0}<(\Omega^{2}+\gamma^{2})/4$. The imaginary part of the energy (i.e., the inverse lifetime), instead, is the same for all modes. Both real and imaginary parts show a bifurcation, which is typical of driven-dissipative non-Hermitian systems [24]. A finite value of the intermode coupling $\lambda_{1}$ creates a non-Hermitian avoided crossing of the frequencies (i.e., bifurcations) close to the momenta $q^{*}$ This emerges as a softening in the imaginary part of the energy around $q^{*}$ (see panel (b) bottom in Fig. 4)
@@
By approaching the critical curves, the imaginary parts of the dispersions further softens until they become zero for some momenta, signaling the emergence of an ordered phase. The softening appears different depending on which side is approached: Upon approaching the second-order phase transition line (panel c bottom in Fig. 4), the softening happens at $q=0$, while approaching the first-order phase transition line (panel b bottom in Fig. 4 ), the finite momentum $q^{*}$ softens. This is remarkable, as this would signal an instability at finite momenta, while we assumed an ordered translation-invariant phase. This suggests that our homogeneous Ansatz for the ordered phase is incomplete, and the inclusion of a spatially-modulated order parameter Ansatz would in fact show the emergence of a phase with broken translational symmetry near the first order transition line. The emergence of spatially patterned phases in driven systems is known and correspond to Faraday waves, which are observed, e.g., in periodically driven condensates, see, e.g., [50; 39; 40]. However, this goes beyond the scope of our work and we will leave it for further investigations.
@@
Excitations in the superconducting phase— We consider now the fluctuations in the ordered superconducting phase. In this case there is always a mode with vanishing imaginary and real part at $q=0$, which corresponds to the Goldstone mode of the ordered phase (panel a, right figure, in Fig. 4). The mode is purely diffusive as $\text{Im}(E(q))\propto-q^{2}$ for small $q$. This is well known for driven-dissipative systems and has been predicted and experimentally verified, e.g., for exciton-polariton systems [64; 8]. Note that such mode is a manifestation of Floquet-Nambu-Goldstone modes discussed in Ref. [45].